OpenStudy (anonymous):

log[2](4)=y how do write this in exponential form? log(base 2) (4)=y

Parth (parthkohli):

\(\Rightarrow \log _24 = y \Longrightarrow 2^y = 4 \)

Parth (parthkohli):

\(\Rightarrow \log_x y = z \Longrightarrow x^z = y \) Using this property.